OpenStudy (anonymous):

4sin^2x-3=0 whats the general solution

OpenStudy (anonymous):

is it a multiple choice ? @lucasbeauchemin

OpenStudy (anonymous):

\[4 \sin ^{2}x-3=0\] \[2\sin ^{2}x=\frac{ 3 }{ 2 }\] \[1-\cos 2x=\frac{ 3 }{ 2 },\cos 2x=-\frac{ 1 }{ 2 }=-\cos \frac{ \pi }{3 }=\cos \left( \pi \pm \frac{ \pi }{ 3} \right)\] \[=\cos \left( 2 n \pi+\pi \pm\frac{ \pi }{ 3 } \right) \] \[2x=2 n \pi+\pi \pm \frac{ \pi }{ 3 },x=\frac{ \left( 2 n+1 \right)\pi }{ 2 }\pm \frac{ \pi }{ 6 }\] where n is an integer

OpenStudy (anonymous):

@lucasbeauchemin sorry mate ..i got it wrong somewhere..the above is right

OpenStudy (anonymous):

i see now thanks